OpenStudy (anonymous):

a right cone has a radius of 3 and a height of 4. what is the slant height of the cone?

OpenStudy (bahrom7893):

s^2=r^2+h^2 s^2=9+16=25 s=5

OpenStudy (anonymous):

thaankyou soo much!